Find the optical density and transmittance of 0.04% -lactose solution, if 1/8 of initial light intensity is absorbed in it.
1
Expert's answer
2020-10-23T07:14:28-0400

By definition, transmittance is:


"T = \dfrac{F_t}{F_i}"

where "F_t" is the radiant flux transmitted by that material and "F_i" is the radiant flux received by that material. If 1/8 of initial light intensity is absorbed in the material, then the transmited ratio is:


"T = \dfrac{F_t}{F_i} = 1-\dfrac18 = \dfrac{7}{8}\approx 0.875"

The optical density (or absorbance) is then:


"D = -\log_{10}T = -\log_{10}0.875 \approx 0.058"

Answer. Transmittance: 0.875. Optical density: 0.058.